In some country’s congress, the number of Representatives is 170 less than 5 times the number of senators. There are a total of 550 members. Find the number of senators and the number of representatives.
Answer by ikleyn(52790)

.
senators + representatives = 550


S        + R               = 550


S        + (5*S -170)      = 550


S + 5S - 170 = 550


6S = 550 + 170 =  720


S = 720/6 = 120.


Answer.  120 senators and  550 - 120 = 430 representatives.

Solved.